OpenStudy (bj11684):

How many pounds are there in 3 and 1 over 4 tons? (1 ton = 2000 pounds)

OpenStudy (anonymous):

\[3\frac{ 1 }{ 4}=\frac{ 3*4+1 }{ 4 }=\frac{ 13 }{ 4 } ~tons\] multiply by 2000 and find the answer.

OpenStudy (anonymous):

6500

OpenStudy (bj11684):

thx

OpenStudy (anonymous):

my pleasure

OpenStudy (bj11684):

I gave you a medal